2010-06-16
==========

Program: CL2 (YSOs, PI: Millan-Gabet et al.)

Observers: Stefan Kraus, Fabien Baron

Telescope: S1-E1-W2
Beams:     B5-B6-B6
POPs:       4- 1- 5

CLASSIC on beam 5&6

beam 5: S1
beam 6: E1
S1 reference cart


LT 20:28 = UT 03:28 = LST 13:08

LT 20:40   *** BASELINE S1E1:
LT 20:40 - Del Sco


YSO: MWC297, VVSer, MWC300 (before UT5), MWC275 (3:36-6:30)
Be:  HD138749 (after 5:17), HD181615 (4:49-8:20), HD149757 (3:36-6:03)

LT 20:40   late with setting up POPs
LT 22:10   problems with CLASSIC
LT 22:50 - HD181615, searching for fringes
           found fringes at +0.63cm (22:53)
LT 23:13 - HD168656 (K=2.8), to find fringes and prepare for V1295Aql
   	   doing NIRO alignment
	   found fringes at -0.803cm (23:48)
LT 23:50 - HD177332 (K=6.2), seeing 6-7cm, early calibrator (about 1hr before target)
   	   recording SYNC 500 Hz (100 Hz Frg), offset -0.869cm (23:58), -0.856 (00:05)
LT 00:05 - HD189509 (K=6.5), early calibrator close to target
   	   E1 has today some problems with acquisition
	   waiting 5 minutes until it comes into delay range (E1 delay goes to 42m not 44m)
	   cannot find fringes at expected position, also FLUOR has problems
	   recording NORM 500 Hz, offset -0.86cm (00:40), interrupting since we want SYNC mode
	   seeing 7-8cm
	   recording SYNC 500 Hz, offset -0.856cm (00:46), more piston than yesterday
LT 01:?? - HD190073
   	   recording SYNC 500 Hz, offset -0.849cm (01:04), -0.8310cm (01:14), clear fringe at 326m BASELINE!
	   lost star on E1 during shutter sequence, re-acquired star during shutter DARK  
	   (might be OK, otherwise use shutter sequence from next data set)
	   therefore repeat data recording with smaller data set
   	   recording SYNC 500 Hz, offset -0.825cm (01:21), seeing 8cm
LT 01:25 - HD189509
   	   lost fringes again, also FLUOR has problem -- this is clearly correlated
	   re-synced clock several times and restarted OPLE gui, which might have fixed the problem
   	   recorded NORM 500 Hz, interrupted
	   recording SYNC 500 Hz, offset -0.7443cm (01:46), -0.735cm (01:49), seeing 8-9cm
	   data looks very good
LT 01:47 - HD190073
	   recording SYNC 500 Hz, offset -0.725cm (02:00), -0.702cm (02:08), -0.69?cm (02:12), counts 120
LT 02:12 - HD189509
   	   recording SYNC 500 Hz, offset -0.615cm (02:24), -0.587cm (02:32), counts 110, seeing 6-7cm
LT 02:32 - HD190073
   	   recording SYNC 500 Hz, offset -0.597cm (02:35), -0.565 (02:44), counts 115
LT 02:50 - HD188385 (other calibrator, used by Eisner+ 2009)
   	   cannot find fringes
LT 03:10 - HD189509
   	   cannot find fringes, seem to be far off, strange thing
   	   recording SYNC 500 Hz, offset -0.327cm (03:28), -0.283cm (03:36), counts 100-110
	   alignment seems to be a bit off
LT 03:35 - HD190327 (K=3.2)
	   redoing alignment
LT 03:40 - HD188385
   	   some problem with S1
	   alignment now better
   	   recording SYNC 500 Hz, offset -0.185cm (03:49), -0.148 (03:56), counts 120
LT 03:56 - HD190073, seeing 8-9cm
   	   recording SYNC 500 Hz, offset -0.155 (04:05), -0.134 (04:08), -0.107 (04:13)
LT 04:17 - HD189509, seeing 7-9
   	   recording SYNC 500 Hz, offset -0.41 (04:20), +0.00 (04:27), counts 110-120
LT 04:30 - HD190073
   	   recording SYNC 500 Hz, scan contains no fringes
	   cannot find fringes
LT 04:46 - HD189509
   	   offset +0.12
LT 04:48 - HD190073
   	   out of delay
LT 04:50 - HD185758
   	   try to get another calibrator
	   recording SYNC 500 Hz, offset +0.28 (05:07)
	   (alignment got worse again)
LT 05:15   closing down